OpenStudy (anonymous):

Write Out The Following Trigonometric Idenetites Pythagorean - There are 3 Cos2x - There are 3 Sin2x I never really learned these (was absent).

OpenStudy (anonymous):

\[\sin ^{2}x+\cos ^{2}x=1\] \[\cos2x=\cos ^{2}x-\sin^{2}x\] \[\sin2x=2sinxcosx\]

OpenStudy (anonymous):

Is that it?

OpenStudy (cwrw238):

others are sec^2 x = 1 + tan^2 x cosec^2 x = 1 + cot^2 x (part 1)

OpenStudy (anonymous):

Part 2?

OpenStudy (cwrw238):

cos 2x = cos^2 x - sin^2 x the other 2 can be found by substituting for cos^2 x and sin^2 x in the above hint: use the first identity that myko posted

OpenStudy (anonymous):

Elaborate a little further, please. I'm missing a lot of this information.

OpenStudy (anonymous):

So far, I've got cos2x = 1-2sin^2 x

OpenStudy (cwrw238):

right thats one of them the other one is obtained by substituting sin^2x = 1 - cos^2 x into the formula for cos 2x

OpenStudy (anonymous):

So would I end up with an equation of cos2x = cos^2x - ( 1 - cos^2 x) ? If so, how would I simplify it? Or is that the answer?

OpenStudy (cwrw238):

cos2x = cos^2x - ( 1 - cos^2 x) = cos^2x - 1 + cos^2 x ( sign changes when parentheses are removed) = 2cos^2 x - 1

OpenStudy (anonymous):

Glad to see I was on the right track. Thanks for the help.

OpenStudy (cwrw238):

yw